# suppSupplementary References

### v1

@@
(June 20, 2024)
We propose a novel instrumental variable to identify fiscal shocks in small open economies. Under the assumptions that unexpected changes in trading partners correlate with output of an open economy and unexpected fiscal shocks of a small economy are unrelated to its trading partners’ forecast errors, we use forecast errors of trading partner economies to proxy unexpected shocks in domestic output. We show that this instrument is relevant and find evidence that supports its exogeneity. Using this IV strategy, we find that the two-year cumulative spending multiplier is around 1 for Canada and 0.5 for euro area small open economies.
@@ 1 Introduction
Estimating the dynamic effects of policy shocks is a central but notoriously hard task in empirical macroeconometrics given the difficulty of credibly identifying exogenous variation in the policy variable (Nakamura and Steinsson, 2018). To overcome this challenge, modern macro-econometrics increasingly uses external instruments to identify dynamic causal effects (Stock and Watson, 2018). One area where the use of these instruments has become prominent is macro-fiscal literature (e.g., Mertens and Ravn, 2013, 2014; Nakamura and Steinsson, 2014; Caldara and Kamps, 2017; Ramey and Zubairy, 2018) that provides estimates on the effects of fiscal policy. [^2] Following the SVAR model of Blanchard and Perotti (2002), much of the prior macro-fiscal research has relied on using assumptions about timing restrictions in the real-world setting of fiscal policy to identify fiscal shocks. Another prominent identification strategy has been to use sign restrictions, as in Mountford and Uhlig (2009). Furthermore, another branch of the literature has utilized narratively identified plausibly exogenous policy changes or other shocks in the identification of the effects of fiscal policy. See, for example, Romer and Romer (2010) and Mertens and Ravn (2013, 2014) in the context of tax policy changes and Ramey (2011) and Ramey and Zubairy (2018) in which a military-news variable is constructed to identify plausibly exogenous variation in government spending. See Ramey (2019) for a comprehensive summary of the macro-fiscal literature. More recently, the importance of statistical identification, which uses heteroskedasticity and higher moments to achieve identification, has grown (see e.g., Lewis, 2021). Given a valid instrument, researchers can estimate structural relationships in the presence of endogenous variables – a situation which is emblematic of macroeconomics.
@@
Our primary contribution to this literature is to introduce a novel instrument for aggregate output shocks that can be used to identify fiscal shocks in a small open economy setting. As shown by Caldara and Kamps (2017), non-fiscal proxy variables that are not correlated with the policy shocks of interest can still be used to uncover these shocks within a SVAR model. Proxies that are exogenous to policy shocks (but correlated with other shocks) solve for the identification problem by allowing the econometrician to estimate the systematic component of fiscal policy. [^3] The systematic component reflects the reaction function of fiscal authorities or, in other words, the implicit fiscal rules that govern how fiscal policy endogenously reacts to other (non-policy) shocks in the model. At a quarterly frequency these reflect, for example, automatic stabilizers that mainly work through output. As fiscal variables may endogenously react to changes in e.g. aggregate output, one needs to account for this reaction in a SVAR to correctly identify the policy shocks from the reduced form residuals. In practice, the elasticities of fiscal variables to other endogenous variables are estimated by instrumenting the relevant non-fiscal variables with proxies that are unrelated to fiscal shocks.
In this paper we utilize the small open economy setting and propose to use professional forecast errors of trading partner economies as a proxy for output shocks. Given that these errors are unrelated to fiscal policy shocks of a small open economy (exogeneity) while at the same time able to explain output variations in the small domestic economy (relevance), the proposed instrument is plausibly valid. We show empirically that trading partner forecast errors are strongly correlated with the unexpected part of domestic output and that the robust first stage F-statistics when applying the instrument exceed the typical threshold value of 10. These results imply that the proposed instrument is not weak.
@@ 3.1 SVAR specification
In lag length selection, we rely on both the standard information criteria and partial autocorrelation functions of the residuals to specify a model that both contains enough information and has no autocorrelation in the residuals. For both the euro area small open economies and Canada, our baseline specification has a lag length of 5.101010In subsection 4.4 and in Online Appendix, we also examine robustness to different choices relating to the specification of the reduced form VAR. For constructing confidence intervals we utilize the residual-based moving block bootstrap proposed by Brüggemann et al. (2016) that is shown to be applicable for Proxy-SVARs / SVAR-IVs. [^11] In this process we also apply Kilian (1998) finite sample correction. Bootstrap confidence intervals are constructed using Efron (1979) percentiles.
@@ 3.3 Data
VAR variables.— We collect data from Statistics Canada, Eurostat and OECD. For $g$, $r$, $gdp$ we rely on national statistical agencies in Statistics Canada and Eurostat while rest of the variables are collected from OECD. We limit ourselves to studying periods with relatively stable macro-institutions, that is, we wish to avoid possible structural breaks. Thus, for Canada, our sample period is 1986Q1-2019Q4 which is motivated by the start of the Great Moderation in the mid 1980s. The sample period for euro area economies is 1999Q1-2019Q4, i.e. the EMU period before COVID-19.161616Due to missing data on some of the variables, the panel is unbalanced. All data are seasonally adjusted when applicable. For Canada we use real government spending but for European economies we deflate government spending and investment by the GDP deflator since deflators for the individual series are not available for all countries of interest.
@@ 5 Conclusion
In this paper, we propose a novel instrument for aggregate output that is based on professional forecast errors in trading partner economies. We apply the instrument in fiscal SVAR-IVs to disentangle policy shocks from endogenous variation in output. The instrument is suited to a small open economy setting where policy shocks of the domestic economy do not cause contemporaneous forecast errors in its trading partners while, at the same time, unexpected developments in these trading partners do cause variation in domestic output.

### v2

@@
(October 3, 2024)
We identify fiscal SVAR-IVs by utilizing unexpected variation in the output of trading partner economies, measured by professional forecast errors, to account for the systematic component of fiscal policy. Our identification builds on the small open economy assumption that these forecast errors correlate with output but are exogenous to domestic fiscal policy. In applying our approach to Canada and euro area small open economies we show that the instrument is relevant and find suggestive evidence for its exogeneity. Our baseline estimates for the two-year cumulative spending multiplier are around 1 for Canada and 0.5 for euro area small open economies.
@@ 1 Introduction
Estimating the dynamic effects of policy shocks is a central but notoriously hard task in empirical macroeconometrics given the difficulty of credibly identifying exogenous variation in the policy variable (Nakamura and Steinsson, 2018). To overcome this challenge, modern macro-econometrics increasingly uses external instruments to identify dynamic causal effects (Stock and Watson, 2018). One area where the use of these instruments has become prominent is macro-fiscal literature (e.g., Mertens and Ravn, 2013, 2014; Nakamura and Steinsson, 2014; Caldara and Kamps, 2017; Ramey and Zubairy, 2018) that provides estimates on the effects of fiscal policy. [^2] Following the SVAR model of Blanchard and Perotti (2002), much of the prior macro-fiscal research has relied on using assumptions about timing restrictions in the real-world setting of fiscal policy to identify fiscal shocks. Another prominent identification strategy has been to use sign restrictions, as in Mountford and Uhlig (2009). Furthermore, another branch of the literature has utilized narratively identified plausibly exogenous policy changes or other shocks in the identification of the effects of fiscal policy. See, for example, Romer and Romer (2010) and Mertens and Ravn (2013, 2014) in the context of tax policy changes and Ramey (2011) and Ramey and Zubairy (2018) in which a military-news variable is constructed to identify plausibly exogenous variation in government spending. See Ramey (2019) for a comprehensive summary of the macro-fiscal literature. More recently, econometric methods using heteroskedasticity and higher moments to achieve identification have gained prominence (see e.g., Lewis, 2021). Given a valid instrument, researchers can estimate structural relationships in the presence of endogenous variables – a situation which is emblematic of macroeconomics.
@@
Our primary contribution to this literature is to introduce an instrumental variable for aggregate output shocks that can be used to identify fiscal shocks in a small open economy setting. As shown by Caldara and Kamps (2017), non-fiscal proxy variables that are not correlated with the policy shocks of interest can still be used to uncover these shocks within a SVAR model. Proxies that are exogenous to policy shocks (but correlated with other shocks) solve for the identification problem by allowing the econometrician to estimate the systematic component of fiscal policy. [^3] The systematic component reflects the reaction function of fiscal authorities or, in other words, the implicit fiscal rules that govern how fiscal policy endogenously reacts to other (non-policy) shocks in the model. At a quarterly frequency these reflect, for example, automatic stabilizers that mainly work through output. As fiscal variables may endogenously react to changes in e.g. aggregate output, one needs to account for this reaction in a SVAR to correctly identify the policy shocks from the reduced form residuals. In practice, the elasticities of fiscal variables to other endogenous variables are estimated by instrumenting the relevant non-fiscal variables with proxies that are unrelated to fiscal shocks.
In this paper we utilize the small open economy setting and propose to use professional forecast errors of trading partner economies as a proxy for output shocks. [^4] To the best of our knowledge, the idea of using unexpected variation in trading partner economies has not been used in the fiscal SVAR-IV context previously. However, following Rigobon (2004), some papers studying the cyclicality of fiscal policy (e.g. Jaimovich and Panizza, 2007; Ilzetzki and Végh, 2008; Vegh and Vuletin, 2015) have used a similar identification strategy in an effort to estimate the endogenous reaction of fiscal policy to changes in output. Unlike our study, which uses professional forecast errors, these papers have used weighted output growth in trading partners outright but have found in some cases this instrument to be weak. Arguably, forecast errors are also less prone to exogeneity concerns given that they should, from the onset, capture unexpected variation in foreign economies. For example, a policymaker aiming to balance business cycles in a small open economy would naturally account for the expected development in foreign economies, thus likely creating a link with fiscal policy. Given that these errors are unrelated to fiscal policy shocks of a small open economy (exogeneity) while at the same time able to explain output variations in the small domestic economy (relevance), the proposed instrument is plausibly valid. We show empirically that trading partner forecast errors are strongly correlated with the unexpected part of domestic output and that the robust first stage F-statistics when applying the instrument exceed the typical threshold value of 10. These results imply that the proposed instrument is not weak.
@@ 3.1 SVAR specification
In lag length selection, we rely on both the standard information criteria and partial autocorrelation functions of the residuals to specify a model that both contains enough information and has no autocorrelation in the residuals. For both the euro area small open economies and Canada, our baseline specification has a lag length of 5.111111In subsection 4.4 and in Online Appendix, we also examine robustness to different choices relating to the specification of the reduced form VAR. For constructing confidence intervals we utilize the residual-based moving block bootstrap proposed by Brüggemann et al. (2016) that is shown to be applicable for Proxy-SVARs / SVAR-IVs. [^12] In this process we also apply Kilian (1998) finite sample correction. Bootstrap confidence intervals are constructed using Efron (1979) percentiles.
@@ 3.3 Data
VAR variables.— We collect data from Statistics Canada, Eurostat and OECD. For $g$, $r$, $gdp$ we rely on national statistical agencies in Statistics Canada and Eurostat while rest of the variables are collected from OECD. We limit ourselves to studying periods with relatively stable macro-institutions, that is, we wish to avoid possible structural breaks. Thus, for Canada, our sample period is 1986Q1-2019Q4 which is motivated by the start of the Great Moderation in the mid 1980s. The sample period for euro area economies is 1999Q1-2019Q4, i.e. the EMU period before COVID-19.171717Due to missing data on some of the variables, the panel is unbalanced. All data are seasonally adjusted when applicable. For Canada we use real government spending but for European economies we deflate government spending and investment by the GDP deflator since deflators for the individual series are not available for all countries of interest.
@@ 5 Conclusion
In this paper, we propose an instrument for aggregate output that is based on professional forecast errors in trading partner economies and apply it in fiscal SVAR-IVs to disentangle policy shocks from endogenous variation in output. The instrument is suited to a small open economy setting where policy shocks of the domestic economy do not cause contemporaneous forecast errors in its trading partners while, at the same time, unexpected developments in these trading partners do cause variation in domestic output.